# On the choice of the non-trainable internal weights in random feature maps

### v1

@@ 7. Summary and future work
We established the notion of good features and good internal parameters for random feature maps with a $\tanh$-activation function. These good internal weights are characterised by affinely mapping the training data into the nonlinear, non-saturated domain of the $\tanh$-activation function. We established that the number of good features $N_{g}\leq D_{r}$ is the controlling factor in determining the forecasting skill of a learned surrogate map, rather than the feature dimension $D_{r}$. Interestingly, the forecasting skill was found to be equally deteriorated by linear features as by non-saturated features. The non-convex set of good features could be written as a union of two convex sets which are mutual reflections of each other. We developed computationally cheap hit-and-run sampling algorithms to uniformly sample from the set of good internal parameters. We demonstrated how ridge regression engages with a given number of good and bad features. It was found that non-saturated features are eliminated almost entirely by the outer weights provided a sufficient number of good features are present. Once there are sufficiently many good features present to allow for a significant reduction of the data mismatch term of the loss function, regularization kicks in and reduces the norm of the outer weights corresponding to good features.
@@
We further showed that a single-layer feedforward network with the same width $D_{r}$ trained with gradient descent exhibits inferior forecasting skill compared to a random feature map surrogate map which used only good internal parameters. The neural network achieves a significantly smaller value of the loss function. Good forecasting skill, however, requires multiple applications of the surrogate map, and as we showed is controlled by the number of good features. The lower forecasting skill is due to the optimization process not finding solutions on the measure-zero set of good parameters. Even when initialized with good parameters, the gradient descent quickly reduces the number of good internal parameters.
The proposed optimization-free algorithm to choose internal non-trainable parameters can potentially lead to new design and computationally cheap training schemes for more complex network architectures. Our algorithms may be used to further improve the forecasting skill of reservoir computing [35, 10]. The distinction into linear, saturated and good features readily translates to other sigmoidal activation functions. It will be interesting to test if larger values of $L_{0,1}$, which we showed in Section 4.3 to lead to a larger effective range explored by features, may improve the forecast skill.
We considered here random feature maps as an alternative to single-layer feedforward networks. It will be interesting to see if the ideas of selecting weights according to the domain of the $\tanh$-activation function can be employed to generate deep random feature networks where at each layer weights and biases are drawn using our hit-and-run algorithm. This is planned for future research.
@@ Acknowledgement
The authors would like to acknowledge support from the Australian Research Council under Grant No. DP220100931. GAG would like to acknowledge discussions with Nicholas Cranch in the early stages of this research.
@@ 8.1. Adaptive learning rate for the single-layer neural network
We tried several other strategies such as finding an optimal learning rate every few steps using bisection, random modifications of the learning rate based on the behavior of the loss function, aggressive constant learning rates and conservative constant learning rates, piecewise linear learning rates etc. We found that the simple strategy presented in Algorithm 3 leads to the lowest final value of the loss function for the same number of gradient descent steps. Figure 20 shows the adaptive learning rate used during the training instance shown in Figure 16.
Figure 20. Adaptive learning rate $\eta$ used during the training instance presented in Figure 16. The associated evolution of the loss function is shown in Figure 16.

### v2

@@ 8. Summary and future work
We established the notion of good features and good internal parameters for random feature maps with a $\tanh$-activation function. These good internal weights are characterised by affinely mapping the training data into the nonlinear, non-saturated domain of the $\tanh$-activation function. We established that the number of good features $N_{g}\leq D_{r}$ is the controlling factor in determining the forecasting skill of a learned surrogate map, rather than the feature dimension $D_{r}$. Interestingly, the forecasting skill was found to be equally deteriorated by linear features as by saturated features. We developed computationally cheap hit-and-run sampling algorithms to uniformly sample from the set of good internal parameters. The hit-and-run Algorithms 1 and 2 sample the internal weights from a data-informed convex set, rather than obtaining them by minimizing some cost function. We demonstrated how ridge regression engages with a given number of good and bad features. In particular, saturated features are eliminated almost entirely by the outer weights, provided a sufficient number of good features are present. Similarly, linear features are suppressed by the outer weights, albeit to a lesser degree. Once there are sufficiently many good features present to allow for a significant reduction of the data mismatch term of the loss function, regularization kicks in and reduces the norm of the outer weights corresponding to good features.
@@
We further showed that a single-layer feedforward network with the same width $D_{r}$ trained with gradient descent exhibits inferior forecasting skill compared to a random feature map surrogate map which used only good internal parameters. The neural network achieves a significantly smaller value of the loss function. Good forecasting skill, however, requires multiple applications of the surrogate map, and, as we showed, is controlled by the number of good features. The lower forecasting skill is due to the optimization process not finding solutions on the measure-zero set of good parameters. Even when initialized with good parameters, the gradient descent quickly reduces the number of good internal parameters.
We found numerically that including global information about the attractor, in form of the constraints for the internal weights (13), is sufficient to ensure that the trained random feature map reproduces the statistical properties of the underlying dynamical system and its invariant measure. The rationale for the choice of good internal weights was entirely motivated by the nonlinear non-saturated structure of the $\tanh$-activation function. How far the constraints on the function domain (13) translate into dynamical information ensuring the preservation of the invariant measure remains an open question, planned for further research.
The proposed optimization-free algorithm to choose internal non-trainable parameters can potentially lead to new design and computationally cheap training schemes for more complex network architectures. Our algorithms may be used to further improve the forecasting skill of reservoir computers [41, 11]. In [37] it is shown that the hit-and-run algorithm can be used for initializing modified RFMs, which include skip connections, and deep architectures, to achieve state-of-the-art forecasting skill, outperforming standard reservoir computers with significantly less computational effort and model sizes.
The parameters $L_{0}$ and $L_{1}$ that define good parameters can, in principle, be considered as hyperparameters, and could be tuned, for example, using Bayesian optimization [26]. We have refrained here from doing so as we have not observed significant changes in the forecasting skill for values close to $L_{0}=0.4$ and $L_{1}=3.5$ which we used throughout this work.
We considered here the $\tanh$-activation function as it is widely used in reservoir computing. The separation of the domain into linear, saturated and good regions can readily be extended to other continuous sigmoidal functions. These activation functions are widely used for random feature maps and in reservoir computing architectures. They are less frequently used for deep neural networks where the saturated regions lead to the vanishing gradient problem [14]. However, we note that recently the $\tanh$-function gained again more interest in the machine learning community [51]. It is an interesting question whether our methodology can be extended to other commonly used activation functions such as the ReLU-type functions or trigonometric functions as initially employed for random feature maps [43]. For example, for GELU [18] one may define good features to be those that correspond to an interval around 00 where the function is nonlinear and non-saturated. For classical Fourier random feature maps, one could separate the domain into linear and saturated regions near the roots and maxima/minima of the $\sin$- and $\cos$-functions, and consider the complement as the “good” region; the linear domain of the $\sin$-function needs to be aligned with the saturated domain of the $\cos$-function, and vice versa. Exploring whether the good regions of such functions allow for sufficient variability of the corresponding features and if this can lead to an improved sampling is planned for future research.
@@ Acknowledgement
The authors would like to acknowledge support from the Australian Research Council under Grant No. DP220100931.
@@ 9.1. Adaptive learning rate for the single-layer neural network
We tried several other strategies such as finding an optimal learning rate every few steps using bisection, random modifications of the learning rate based on the behavior of the loss function, aggressive constant learning rates, conservative constant learning rates, piecewise linear learning rates etc. We found that the simple strategy presented in Algorithm 3 leads to the lowest final value of the loss function for the same number of gradient descent steps. Figure 23 shows the adaptive learning rate used during the training instance shown in Figure 17.
Figure 23. Adaptive learning rate $\eta$ used during the training instance presented in Figure 17, where the loss function is shown.